OpenStudy (anitas.):

What is the value of n in the equation 3n + 2(n + 2) = 9n + 12? I need help explanations step by step.

alones (alones):

This thing ill explain it beettte :P http://mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

\[ 3n + 2(n + 2) = 9n + 12? \] step by step right?

OpenStudy (anonymous):

ready?

OpenStudy (anitas.):

yes

OpenStudy (anonymous):

distribute the 2 first for the \(2(n+2)\) part what do you get/

OpenStudy (anitas.):

4?

OpenStudy (anonymous):

oh no lets go slow

OpenStudy (anonymous):

your answer has to have an \(n\) in it \[2(n+2)=2\times n+2\times 2\] by the distributive law

OpenStudy (anonymous):

we would write that as \[2(n+2)=2n+4\]

OpenStudy (anonymous):

start with \[3n + 2(n + 2) = 9n + 12\] next we have \[3n+2n+4=9n+12\] next step is to combine like terms on the left

OpenStudy (anitas.):

5n+4=9n+12

OpenStudy (anitas.):

and then i subtract 5n from both sides right?

OpenStudy (anonymous):

yes

OpenStudy (anitas.):

ok would that be 4=9n+12−5n?

OpenStudy (anonymous):

yes, but you probably want to subtract \(9n-5n\) instead of writing it like you did

OpenStudy (anonymous):

what is \(9n-5n\)?

OpenStudy (anitas.):

4n

OpenStudy (anonymous):

right so \[4=4n+12\]

OpenStudy (anonymous):

next step?

OpenStudy (anitas.):

8+12n

OpenStudy (anonymous):

ok no

OpenStudy (anonymous):

i said that wrong the n in on the RIGHT you want the number on the LEFT subtract 12 from both sides

OpenStudy (anonymous):

\[4=4n+12\] subtract \(12\) what do you get?

OpenStudy (anitas.):

-8=4n right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now divide both sides by \(4\) to get \(n\) by itself

OpenStudy (anitas.):

-8/44=n

OpenStudy (anitas.):

oops i mean -8.4=n

OpenStudy (anonymous):

not 44, just 4 \[\frac{-8}{4}=n\]

OpenStudy (anonymous):

what is minus 8 divided by 4?

OpenStudy (anitas.):

lol sorry i am gettin this wrong i ment t write that lol sorry

OpenStudy (anonymous):

what number is it?

OpenStudy (anitas.):

-2

OpenStudy (anonymous):

bingo

OpenStudy (anitas.):

ok thank you very much!! <33333 btw why did ur ss go from 100 to 99?

OpenStudy (anonymous):

i got hacked apparently

OpenStudy (anitas.):

oh dats sad. i noticed though

OpenStudy (anonymous):

i will get it back